Water is falling from a height of 35m, Assuming that 3/4 of the K.E of the falling water gets converted to heat,then what is the rise in temperature?
1
Expert's answer
2013-11-27T10:17:01-0500
Kinetic energy equals change of potential energy:

E = mgh

Changes in temperature can be found as

Q=c*m*t

t=Q/(c*m)

where c is specific heat capacity, c=4200 J/(kg*K). Recalling that Q=3/4 E, we find:
t=3/4 mgh/ (c*m) = 3*9.8*35/(4*4200) = 0.06125 K